OpenStudy (anonymous):

Find the value of the discriminant for the following quadratic equation: 2x^2 + 6x + 5 = 0

OpenStudy (sidsiddhartha):

b^2-4*a*c=discriminant for ax^2+bx+c

OpenStudy (whpalmer4):

To be picky, the discriminant of a quadratic (sometimes written \(\Delta\)) is defined as \[\Delta=b^2-4ac\]only if the values of \(a,b,c\) come from \[ax^2+bx+c=0\] \[ax^2+bx+c=3\] does not have the above as a discriminant (must have all non-zero terms on same side before computing discriminant) and \[ax^2+bx+c\]does not have a discriminant, not being an equation.